How much tension must a cable withstand if it used to accelerate a 1500 kg object vertically upward at 0.80 m/s²?
1
Expert's answer
2020-10-05T10:55:49-0400

Let's apply the Newton's Second Law of Motion:


"\sum F_y = ma_y,""T - mg = ma,"

here, "T" is the tension in the cable, "mg" is the weight of the object, "m = 1500 \ kg" is the mass of the object, "a = 0.80 \ \dfrac{m}{s^2}" is the acceleration of the object and "g = 9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity.

Then from this equation we can find the tension in the cable:


"T = m(g+a),""T = 1500 \ kg \cdot (9.8 \ \dfrac{m}{s^2}+0.80 \ \dfrac{m}{s^2}) = 15900 \ N."

Answer:

"T = 15900 \ N."